# SCINE — Software for Chemical Interaction Networks

### v1

@@ II.9.1 Field of application
Puffin is a program that is launched in high numbers on high-performance computing infrastructure in order carry out composite workflows. The program is currently employed in, but not limited to, our exploration framework as it can encode any algorithm that requires no more than a set of atomistic input structures and an electronic structure model. Therefore, possible fields of application are high-throughput campaigns that carry out individual energy evaluations, optimization algorithms, or more complex workflows based on chained optimization algorithms such as reaction trials that combine reactive complex formation, Newton trajectory, single-ended transition state search, intrinsic reaction coordinate calculations, and additional checks to assess the bonding patterns, molecular charge and spin multiplicity of any found product(s).
To carry out these workflows, Puffin depends on other SCINE modules such as ReaDuct and possibly external quantum chemistry software. To simplify the setup of this large software stack, we built a so-called bootstrapping procedure through which a Puffin instance automatically downloads and compiles all required software. Large-scale deployments on high-throughput computing infrastructure are further simplified by running Puffin in a virtualized environment. For this, we currently support Docker and Apptainer images.
@@ II.10.1 Field of application
In case the exploration is carried out autonomously, i.e., calculations are setup iteratively based on the already explored reaction space, it is desirable to employ the same database for both, carrying out calculations and storing the CRN. However, this requires the database to be able to handle thousands of simultaneous queries from the algorithms driving the exploration and the processes that carry out individual electronic structure calculations. Hence, we have developed a database schema for a MongoDBMongoDB Inc. database and a custom wrapper to interact with the databaseBensberg et al. (2023g). The wrapper ensures that every interaction with the database is synchronized with its current state. This guarantees that multiple processes can act on the current database simultaneously.
@@
We refer to Ref. 178 for more details on the individual definitions. The relations between all documents in the SCINE Database are illustrated in Fig. 4.
@@ II.10.2 API
In order to retrieve information from the database, the collections objects hold the querying methods query_x, iterate_x, and count, x being a placeholder for the applied collection such as query_calculations or query_structures. The querying methods (cf. Listing 10) can take any valid MongoDB query, which are given as documents that specify which keys of the database documents should fulfill certain criteria, e.g., equality of values. More complex matches are achieved with operators, which start with \$, and multiple queries can be combined with logical operators such as “and”/“or”.
@@
The underlying electronic structure methods can generally be summarized in the Model objects, which can be directly queried for. The results of the queries can be built directly into loops to aggregate information. For example, one can collect the Cartesian coordinates of all structures that were optimized with GFN2-xTB, but have already received a single-point energy obtained with an exchange-correlation density functional:
@@ II.11.2 API
A steered exploration can be carried out in a Python environment by constructing the linear exploration protocol. The example in Listing 13 sets up an exploration that adds two compounds to the database, creates conformers, selects the lowest energy conformers in each cluster based on k-means clustering, and samples possible reactions between the compounds. Later, the protocol is enhanced with a selection of all found products with updated reactive site rules and then only unimolecular rearrangements are sampled. Due to the linear execution logic of the Steering Wheel, it does not matter whether the two last steps were added later or from the beginning of the exploration run as both procedures produce identical results.
@@ 35    Association(
Due to the top-level approach and flexibility of the Steering Wheel, the decision on the exploration protocol requiresva priori either a hypothesis about the possible reaction mechanism or a good overview of the status of the exploration in order to adapt the exploration strategy on the fly. Because the first requirement is rare, we have built the framework around the second requirement by integrating it into our graphical user interface Heron. This allows one to access the current exploration status and preview the effect of potential further exploration steps.
@@ II.12.1 Field of application
Microkinetic modeling simulations are crucial to connect the microscopic description of a chemical reaction as a detailed reaction network to macroscopic observations such as yields by predicting the concentration fluxes through the network. Microkinetic modeling simulations can be run through the SCINE module KiNetXBensberg et al. (2023d). Moreover, such simulations can automatically steer the reaction network explorationBensberg and Reiher (2023a) by interweaving them directly in a rolling exploration. KiNetX is interfaced with the library SundialsHindmarsh et al. (2005); Gardner et al. (2022), which provides efficient integration routines for the differential equations in the microkinetic modeling. As an alternative to KiNetX, SCINE provides an interface to the program Reaction Mechanism Simulator (RMS)Johnson, McGill, and Green (2022); Johnson et al. (2023) for microkinetic modeling. To predict which reactions and compounds in a reaction network dominate the overall kinetics and therefore must be described accurately, the SCINE framework provides sensitivity analysis approaches for the RMS microkinetic modeling through the library SALibHerman and Usher (2017); Iwanaga, Usher, and Herman (2022). The sensitivity analysis can also be interwoven with a rolling exploration to identify the key reactions and compounds and refine their descriptions by reoptimizing structures or recalculating electronic energies with more accurate electronic structure modelsBensberg and Reiher (2023c).
@@ II.12.3 API
As a result, the integrate function returns the final and maximum concentrations, the integrated absolute concentration fluxesBensberg and Reiher (2023a) through the compounds and reactions.
III Conclusions
@@ Acknowledgements.
This publication was created as part of NCCR Catalysis, a National Centre of Competence in Research funded by the Swiss National Science Foundation (grant number 180544). M. E. gratefully acknowledges an ETH Zurich Postdoctoral Fellowship. M. S. gratefully acknowledges a Swiss Government Excellence Scholarship for Foreign Scholars and Artists. M. M. gratefully acknowledges financial support through ETH Research Grant ETH-43 20-2.
@@
Data sharing is not applicable to this article as no new data were created or analyzed in this study.

### v3

@@ II.9.1 Field of application
Puffin’s main aim is to combine basic algorithms implemented in dedicated electronic structure packages or abstracted by other SCINE modules into larger tasks representing a single calculation in network exploration. These calculations can ultimately be simple tasks, such as calculating a single-point energy for a given structure or, as seen in Figure 6, they can also be rather involved. Figure 6 shows the diagram of a single elementary-step trial, including fall-back strategies and alternative branches based on intermediate results. Multiple programs and a direct connection to the exploration database are used. Those steps marked to be carried out by ReadDuct can use any electronic structure program interfaced as a SCINE Core Calculator.
Figure 6: Flowchart of an elementary-step search as implemented in SCINE Puffin.
Puffin is a program that is launched in high numbers on high-performance computing infrastructure in order carry out composite workflows. At startup, the user can assign the computational resources (such as the number of processors and the amount of memory) which a Puffin instance is allowed to utilize. These resource limits will subsequently be propagated by Puffin to any further program it invokes. SCINE Puffin is currently employed in, but not limited to, our exploration framework as it can encode any algorithm that requires no more than a set of atomistic input structures and an electronic structure model. Therefore, possible fields of application are high-throughput campaigns that carry out individual energy evaluations, optimization algorithms, or more complex workflows based on chained optimization algorithms such as reaction trials that combine reactive complex formation, Newton trajectory, single-ended transition state search, intrinsic reaction coordinate calculations, and additional checks to assess the bonding patterns, molecular charge and spin multiplicity of any found product(s).
To carry out these workflows, Puffin depends on other SCINE modules such as ReaDuct and possibly external quantum chemistry software. To simplify the setup of this large software stack, we built a so-called bootstrapping procedure through which a Puffin instance automatically downloads and compiles all required software. With this bootstrap procedure, installing the entire SCINE toolchain becomes straightforward even for unexperienced users. Large-scale deployments on high-throughput computing infrastructure are further simplified by running Puffin in a virtualized environment. For this, we currently support Docker and Apptainer images.
@@ II.10.1 Field of application
In case the exploration is carried out autonomously, i.e., calculations are set up iteratively based on the already explored reaction space, it is desirable to employ the same database for both, carrying out calculations and storing the CRN. However, this requires the database to be able to handle thousands of simultaneous queries from the algorithms driving the exploration and the processes that carry out individual electronic structure calculations. Hence, we have developed a database schema for a MongoDBMongoDB Inc. (2023) database and a custom wrapper to interact with the databaseBensberg et al. (2023g). This wrapper provides useful functionality geared towards the exploration of CRNs which is not available from MongoDB itself. For example, entities such as reactions and structures are defined by the wrapper, as well as related functions (such as retrieving the atoms constituting a given structure). The wrapper has been designed such that every interaction with the database is synchronized with its current state. This guarantees that multiple processes can act on the current database simultaneously.
@@
We refer to Ref. 200 for more details on the individual definitions. The relations between all documents in the SCINE Database are illustrated in Fig. 7.
@@ II.10.2 API
In order to retrieve information from the database, the collections objects hold the querying methods query_x, iterate_x, and count, x being a placeholder for the applied collection such as query_calculations or query_structures. The querying methods (cf., Listing 10) can take any valid MongoDB query, which are given as documents that specify which keys of the database documents should fulfill certain criteria, e.g., equality of values. More complex matches are achieved with operators, which start with \$, and multiple queries can be combined with logical operators such as “and”/“or”.
@@
The underlying electronic structure methods can generally be summarized in the Model objects, which can be directly queried for. The results of the queries can be built directly into loops to aggregate information. For example, one can collect the Cartesian coordinates of all structures that were optimized with GFN2-xTB, but have already received a single-point energy obtained with an exchange–correlation density functional:
@@ II.11.3 API
A steered exploration can be carried out in a Python environment by constructing the linear exploration protocol. The example in Listing 13 sets up an exploration that adds two compounds to the database, creates conformers, selects the lowest energy conformers in each cluster based on k-means clustering, and samples possible reactions between the compounds. Due to the linear execution logic of the Steering Wheel, it does not matter whether the two last steps were added later or from the beginning of the exploration run as both procedures produce identical results.
@@ 35 Association(
Due to the top-level approach and flexibility of the Steering Wheel, the decision on the exploration protocol requires a priori either a hypothesis about the possible reaction mechanism or a good overview of the status of the exploration in order to adapt the exploration strategy on the fly. Because the first requirement is rare, we have built the framework around the second requirement by integrating it into our graphical user interface Heron. This allows one to access the current exploration status and preview the effect of potential further exploration steps.
@@ II.12.1 Field of application
Microkinetic modeling simulations are crucial to connect the microscopic description of a chemical reaction as a detailed reaction network to macroscopic observations such as yields by predicting the concentration fluxes through the network. Microkinetic modeling simulations can be run through the SCINE module KiNetXBensberg et al. (2023d). The version of KiNetX available in SCINE is written in modern C++ and is based on its original MATLAB implementationProppe and Reiher (2019, 2022). Microkinetic modeling simulations can automatically steer the reaction network explorationBensberg and Reiher (2023b) by interweaving them directly in a rolling exploration. KiNetX is interfaced with the library SundialsHindmarsh et al. (2005); Gardner et al. (2022), which provides efficient integration routines for the differential equations in the microkinetic modeling. As an alternative to KiNetX, SCINE provides an interface to the program Reaction Mechanism Simulator (RMS)Johnson, McGill, and Green (2022); Johnson et al. (2023) for microkinetic modeling. To predict which reactions and compounds in a reaction network dominate the overall kinetics and therefore must be described accurately, the SCINE framework provides sensitivity analysis approaches for the RMS microkinetic modeling through the library SALibHerman and Usher (2017); Iwanaga, Usher, and Herman (2022). The sensitivity analysis can also be interwoven with a rolling exploration to identify the key reactions and compounds and refine their descriptions by reoptimizing structures or recalculating electronic energies with more accurate electronic structure modelsBensberg and Reiher (2023a).
@@ II.12.3 API
As a result, the integrate function returns the final and maximum concentrations, the integrated absolute concentration fluxesBensberg and Reiher (2023b) through the compounds and reactions.
III Example Application
To illustrate how individual SCINE modules can be combined to perform a complex task, we study the reaction between ferrocene carboxaldehyde and (S)-alanine (cf., Fig. 9). This reaction is the first step of the synthesis of (R)-$\alpha$-methyl phenylalanine presented by Alonso and Davies in 1995Alonso and Davies (1995). Note that the original reaction sequence involves the sodium salt of (S)-alanine rather than (S)-alanine.
Figure 9: Example reaction studied in this work.
All scripts used to set up and carry out this exploration as well as all data collected are available on ZenodoWeymuth et al. (2024). For the sake of simplicity and speed, we conducted the exploration with the GFN2-xTB methodBannwarth, Ehlert, and Grimme (2019). Furthermore, we did not consider the conformational space of the individual molecules, instead starting from predefined conformations for the two reactants.
To begin the exploration, the code presented in Listing 16 establishes a connection to a database, and stores the structures contained in the files ferrocene_carboxaldehyde.xyz and alanine.xyz as guess structures in the database. They are then subsequently optimized with the GFN2-xTB method (as defined by the model); these optimizations are done by a Puffin instance, which itself relies on ReaDuct to carry out the actual optimization (ReaDuct, in turn, will rely on an optimizing algorithm implemented in the Utilities package, and the external xtb program to calculate electronic energies, nuclear gradients, and Hessian matrices).
The compound_engine will automatically analyze the optimized structures, requesting bond orders to be calculated. This information is then processed by Molassembler to create a molecular graph, which is utilized to assign each structure to a compound as explained above. The thermo_engine automatically requests Hessian calculations for any structures (if such calculations do not yet exist) for later thermochemical analyses. In our minimal example, we use a very simple kinetics_gear which activates any compound for further exploration. In a more sophisticated approach, one could instead employ a different gear relying on KiNetX to enable the exploration of compounds based on their concentration fluxes (see above).
The reaction_engine assigns any elementary steps to the correct reaction. At this stage of the exploration, this engine does not alter the database, since no elementary steps are yet present. Finally, the scheduling_engine takes care of scheduling the individual calculations.
1# Connect to database
2manager = db.Manager()
3db_name = "scine_demonstration"
4ip = ’127.0.0.1’
5port = ’27017’
6credentials = db.Credentials(ip, int(port), db_name)
7manager.set_credentials(credentials)
8manager.connect()
9
10# Define model for all calculations
11model = db.Model("gfn2", "gfn2", "")
12
13# Initialize database
14wipe = True
15if wipe:
16 manager.wipe()
17 manager.init()
18 time.sleep(1.0)
19 # Load initial data
20 ferrocene_carboxaldehyde = "ferrocene_carboxaldehyde.xyz"
21 alanine = "alanine.xyz"
22
23 structure, calculation = insert_initial_structure(
24 manager,
25 ferrocene_carboxaldehyde,
26 0,
27 1,
28 model,
29 settings=default_opt_settings()
30 )
31 structure, calculation = insert_initial_structure(
32 manager,
33 alanine,
34 0,
35 1,
36 model,
37 settings=default_opt_settings()
38 )
39
40# Compound generation and sorting
41compound_gear = BasicAggregateHousekeeping()
42compound_gear.options.model = model
43compound_engine = Engine(credentials)
44compound_engine.set_gear(compound_gear)
45compound_engine.run()
46
47# Thermo-chemical data completion
48thermo_gear = BasicThermoDataCompletion()
49thermo_gear.options.model = model
50thermo_gear.options.job = db.Job("scine_hessian")
51thermo_engine = Engine(credentials)
52thermo_engine.set_gear(thermo_gear)
53thermo_engine.run()
54
55# Sorting elementary steps into reactions
56reaction_gear = BasicReactionHousekeeping()
57reaction_engine = Engine(credentials)
58reaction_engine.set_gear(reaction_gear)
59reaction_engine.run()
60
61# Driving exploration based on kinetics
62kinetics_gear = MinimalConnectivityKinetics() # activate all compounds
63kinetics_engine = Engine(credentials)
64kinetics_engine.set_gear(kinetics_gear)
65kinetics_engine.run()
66
67# Calculation scheduling
68scheduling_gear = Scheduler()
69scheduling_gear.options.job_counts = {
70 "scine_single_point": 500,
71 "scine_geometry_optimization": 500,
72 "scine_ts_optimization": 500,
73 "scine_bond_orders": 500,
74 "scine_hessian": 200,
75 "scine_react_complex_nt2": 100,
76 "scine_dissociation_cut": 100,
77 "conformers": 20,
78 "final_conformer_deduplication": 20,
79 "graph": 1000,
80}
81scheduling_engine = Engine(credentials)
82scheduling_engine.set_gear(scheduling_gear)
83scheduling_engine.run()
Listing 16: Setting up the database for the example exploration. For the sake of brevity, only parts of the full initialization script are shown here. The full script is available on ZenodoWeymuth et al. (2024).
After having initialized the database with the two reactants, we can set up reaction trials in a next step. A brute-force approach would result in over 2’000’000 trial calculations; however, we can cut down this number significantly by taking advantage of chemical intuition as shown in Listing 17. First, we can focus exclusively on bimolecular reactions between (S)-alanine and ferrocene carboxaldehyde, leaving away any unimolecular reactions with the setting enable_unimolecular_trials = False. Furthermore, we employ the predefined CatalystFilter to restrict bimolecular reaction trials to involve ferrocene carboxaldehyde and (S)-alanine, excluding trials with two identical molecules (as the name suggests, this filter is particularly useful in catalysis applications, but it can be very handy also in other situations, such as the one we are dealing with here). Finally we can restrict the exploration to selected functional groups with the variable reactive_site_filter. In our example, we restrict the reactivity to the aldehyde and the amino groups.
1nt_job = db.Job("scine_react_complex_nt2")
2nt_settings = default_nt_settings()
3
4elementary_step_gear = BruteForceElementarySteps()
5elementary_step_gear.trial_generator = BondBased()
6elementary_step_gear.trial_generator.options.model = model
7elementary_step_gear.options.enable_bimolecular_trials = True
8elementary_step_gear.options.enable_unimolecular_trials = False
9
10# Add filters
11elementary_step_gear.trial_generator.reactive_site_filter = AtomRuleBasedFilter(DistanceRuleSet({
12 ’C’: FunctionalGroupRule(0, ’C’, (3, 3), {’O’: 1, ’H’: 1, ’C’: 1}, True),
13 ’O’: DistanceRuleAndArray([FunctionalGroupRule(0, ’O’, (1, 1), {’C’: 1}, True), FunctionalGroupRule(1, ’C’, (3, 3), {’O’: 1, ’C’: 1, ’H’: 1}, True)]),
14 ’N’: FunctionalGroupRule(0, ’N’, (3, 3), {’C’: 1, ’H’: 2}, True),
15 ’H’: FunctionalGroupRule(1, ’N’, (3, 3), {’H’: 2, ’C’: 1}, True)}))
16
17elementary_step_gear.aggregate_filter = CatalystFilter({’Fe’: 1})
18
19# Run
20elementary_step_engine = Engine(credentials)
21elementary_step_engine.set_gear(elementary_step_gear)
22elementary_step_engine.run()
Listing 17: Setting up reaction trials between alanine and ferrocene carboxaldehyde. For the sake of brevity, only parts of the full initialization script are shown here. The full script is available on ZenodoWeymuth et al. (2024).
These reaction trials are subsequently evaluated by Puffin as depicted in Fig. 6. Several elementary steps are found in which the nitrogen atom attacks the carbonyl carbon atom of ferrocene carboxaldehyde, resulting in intermediates 4a and 4b in Fig. 10. Note that before this attack, (S)-alanine coordinates to the iron atom (intermediate 3 in Fig. 10) — Chemoton finds this reaction even though the metal center was not included as a reactive site when setting up the reaction trials.
Figure 10: Energy diagram of the example reaction.
Starting from intermediates 4a and 4b, new reaction trials were created with slightly different filter settings (the script to create these additional trials is also available on ZenodoWeymuth et al. (2024)). These result in several elementary steps eliminating water, leading to intermediates 5a and 5b (cf., Fig. 10). Directly after the formation of water, it forms a hydrogen bond to the nitrogen atom; breaking this hydrogen bond yields the final product (6a and 6b in Fig. 10). Note that Chemoton does not only find the product 6b as shown in Fig. 9, but also the compound 6a in which the carboxylic acid coordinates to the iron atom. This necessitates a deformation of the ferrocene moiety, significantly tilting one of the cyclopentadienyl rings, which results in 6a being less stable than 6b.
IV Conclusions
@@ Acknowledgements.
This publication was created as part of NCCR Catalysis, a National Centre of Competence in Research funded by the Swiss National Science Foundation (grant number 180544). M. E. has been supported by an ETH Zurich Postdoctoral Fellowship and M. S. by a Swiss Government Excellence Scholarship for Foreign Scholars and Artists. We gratefully acknowledge financial support for M. M. through ETH Research Grant ETH-43 20-2. We would like to thank Christoph Brunken for his foundational work on the Swoose module. Moreover, we are grateful to other members of the Reiher research group, students, and collaborators who have contributed to parts of the software (see the Zenodo references cited in this work). In particular, we would like to mention Alberto Baiardi, Robin Feldmann, Nina Glaser, Stefan Gugler, Moritz Haag, Michael Heuer, Tamara Husch, Lucas Lang, Severin Polonius, Jonny Proppe, Gregor Simm, Christopher Stein, and Alain Vaucher.
@@
The data that support the findings of this study are openly available at Zenodo at http://doi.org/10.5281/zenodo.11048378. The source code of the SCINE software modules is available on https://github.com/qcscine as well as on ZenodoBensberg et al. (2023a, b, c); Brunken, Csizi, and Reiher (2021); Bensberg et al. (2023d, e); Bosia et al. (2023a); Bensberg et al. (2023f, 2022); Bosia et al. (2023c); Baiardi et al. (2023); Bensberg et al. (2023g). Additional information about the SCINE modules can be obtained on https://scine.ethz.ch.